OpenStudy (anonymous):

A particle moves along with a velocity of v(t) = 2t + 3 Find the total distance traveled from t=2 to t=5

OpenStudy (ybarrap):

$$ \large \int_2^5v(t)dt\\ \int_2^52t+3~dt\\ =\left (\cfrac{2t^2}{2}+3t\right)\Huge{|^5_2}\\ =(25+15)-(4+6)=?~ $$

OpenStudy (ybarrap):

Do you know why this works?

OpenStudy (anonymous):

so its 30? i think so.

OpenStudy (ybarrap):

It works because, velocity = distance traveled / time taken. So, $$ \int v(t)~dt=\int \cfrac{dx}{dt}dt=\int dx $$ Which is just summing the infinitesimal distances, \(dx\) from 2 to 5.

OpenStudy (anonymous):

thanks!

OpenStudy (ybarrap):

you're welcome